OpenStudy (anonymous):

Find the equation of the line through the point (-2,3) and perpendicular to the line 2x+3y=12. Write the final answer in slope intercept form.

OpenStudy (anonymous):

so 2x+3y = 12, 3y = -2x + 12, y = -2/3 x + 4. So now what is the slope of this line?

OpenStudy (anonymous):

remember: y = mx +b where: m is the slope. b is the y-intercept

OpenStudy (anonymous):

The slope is now -2/3, right?

OpenStudy (anonymous):

Yes. the slope is -2/3. Hmm.. so you have any idea of what is the product of slopes of perpendicular lines?

OpenStudy (anonymous):

I honestly don't know where to go from here...

OpenStudy (anonymous):

Ok there is a concept here. 

OpenStudy (anonymous):

"The products of slopes of perpendicular lines equals -1."

OpenStudy (anonymous):

Is -2/3x+5/3 correct?

sam (.sam.):

Or you can use \[m_1m_2=-1\] Is this familiar to you?

sam (.sam.):

\[m_1=current~slope \\ \\ m_2=perpendicular~slope\]

OpenStudy (texaschic101):

isn't the slope, since it is a perpendicular line, = -3/2 ?

sam (.sam.):

So, \[(-\frac{2}{3})(m_2)=-1 \\ \\ m_2=\frac{3}{2}\]

sam (.sam.):

Form a new equation using the point-slope form with \(m_2\) and coordinate (-2,3) \[y-y_1=m_2(x-x_1)\]

sam (.sam.):

\[y-3=\frac{3}{2}(x-(-2)) \\ \\ y=\frac{3}{2}x+3+3 \\ \\ y=\frac{3}{2}x+6\]

sam (.sam.):

Do you get it? @ebeverly

OpenStudy (anonymous):

I'll have to rework the problem. I got a completely different answer, but this is super helpful. Thank you very much!

sam (.sam.):

Alright, yw :)